OpenStudy (anonymous):

How do I solve these types of problems? 'A ferry heads west across a river. It travels 20 miles per hour in still water. The river flows south at 15 miles per hour. In which direction will the ferry actually travel. How fast will it travel?' I need someone to explain this, so that I can do the other problems like it.

OpenStudy (amistre64):

its a vector related problem

OpenStudy (amistre64):

add the vector to get the red travel distance vector

OpenStudy (amistre64):

<20,0> is the westward vector ; and <0,-15> is the southward vector

OpenStudy (amistre64):

the new vector is simply: <20,0> <0,-15> -------- <20,-15>

OpenStudy (amistre64):

the direction traveled is the tan^-1 (y/x) + pi

OpenStudy (amistre64):

-15/20 = -3/4 tan^-1(-3/4) =abt 233.13 degrees if i did it right

OpenStudy (amistre64):

had to add 270 to it to get itin the proper qudrant

OpenStudy (amistre64):

and the distance traveled is just the pythag of the y and x components

OpenStudy (amistre64):

any of this make sense?

OpenStudy (anonymous):

No...I wish though.

OpenStudy (amistre64):

hmm.... a vector is just an arrow that points in any direction and indicates the distance traveled; they can be called directed line segments

OpenStudy (amistre64):

when we add them, they form a new arrow that measures a new direction and length

OpenStudy (amistre64):

you essentially have the ferry being thrown off course by the river; by 15 meters every sec right?

OpenStudy (amistre64):

well; 15 miles for every hour travled

OpenStudy (amistre64):

if we pythag the new direction; we get 20 west; and 15 south; those are the legs of a triangle and the distance actually covered in one hour is the hypotenuse of that triangle

OpenStudy (amistre64):

sqrt(20^2 + 15^2) = distance traveled in one hour

OpenStudy (amistre64):

in this case; its moved 25 miles in one hour...... or 25 mph

OpenStudy (amistre64):

the direction it moved is to the south and west...measured from some starting point; usually idicated as due est

OpenStudy (amistre64):

due east is the inital starting point of angle measures...

OpenStudy (amistre64):

any of this ringing bells yet :)

OpenStudy (anonymous):

I understand better now, thank you for helping me so much :)